OpenStudy (anonymous):

The product of two positive numbers is 70. Find the numbers if their sum is a minimum.

OpenStudy (anonymous):

you want x+y to be min when x*y=70 so x=70/y x+y=70/y+y (you need the minimum of this) you can do this by taking the derivative of it and making it equal to 0 -70/y^2+1=0 y^2=70 y=+-sqrt70 as it is a positive number the answer is sqrt70

OpenStudy (anonymous):

both numbers are sqrt 70

jhonyy9 (jhonyy9):

35 and 2

OpenStudy (anonymous):

35+2=37 for example 10+7=17 so clearly 35 and 2 cannot be the minimum